What is the sum or difference in 4/5 and 7/8


1
Expert's answer
2020-10-01T15:16:21-0400

"\frac{4}{5}+\frac{7}{8}=\frac{4•8+7•5}{5•8}=\frac{32+35}{40}=\frac{67}{40}"

"\frac{4}{5}-\frac{7}{8}=\frac{4•8-7•5}{5•8}=\frac{32-35}{40}=-\frac{3}{40}".